# A Triadic Suffix Tokenization Scheme for Numerical Reasoning

### 17 Apr 2026 v2

@@ 3.4 Vocabulary Considerations
  • Option A (Separate tokens): Keep digit groups and suffixes as separate tokens. Vocabulary adds only 10 new tokens (k, m, b, t, q, p, pp, ppp, pppp, ppppp). This minimally expands the vocabulary; models may learn to combine digits with suffixes.
@@
Option B produces shorter input sequences (lower token count) and presents the model with ready‑made magnitude‑digit units, eliminating any ambiguity about which suffix belongs to which digit group. The smaller vocabulary of Option A is appealing, but the trade-off between sequence length and vocabulary size remains an empirical question.
While we focus on $N=3$ (Triadic) for its alignment with human readability and its effective balance between vocabulary size and sequence length, the framework is fundamentally designed for any group size $N$, allowing for precise optimization of the model’s vocabulary footprint. It is worth noting that Option B produces the exact same number of input tokens as a simple division into groups of three. Even in the extreme case of $N=1$, the input context length remains no larger than that of standard digit-by-digit representations, as each digit is fused with its magnitude marker into a single atomic token. While Option A is more ”token-greedy” because it emits separate markers, it may still be practical for comparatively larger group sizes ($N\geq 3$) where the relative overhead becomes negligible.
@@ 6.2 Limitations and Future Work
  • Alternative grouping sizes: The choice of 3-digit groups (base‑1000) is natural for human readability, but other sizes could be considered. Language-specific grouping (e.g., 3-2-2 in India, 4-4 in China) can be handled by the parser without changing the core vocabulary [12]. The decision should be guided by what representation is most efficient for the model.

### 20 Apr 2026 v3

@@ 3.4 Vocabulary Considerations
  • Option A (Separate tokens): Keep digit groups and suffixes as separate tokens. Vocabulary adds only 10 new tokens (k, m, b, t, q, p, pp, ppp, pppp, ppppp). This minimally expands the vocabulary; models learn to combine digits with suffixes through attention.
@@
Option B produces shorter input sequences and presents the model with ready-made magnitude-digit units, eliminating ambiguity about which suffix belongs to which digit group. While the smaller vocabulary of Option A is appealing, the trade-off remains an empirical question.
While we focus on $N=3$ (Triadic) for its alignment with human readability and its effective balance between vocabulary size and sequence length, the framework is fundamentally designed for any group size $N$, allowing for precise optimization of the model’s vocabulary footprint. Generally, Option A is more practical for larger group sizes ($N\geq 3$) where the relative token overhead is negligible.
As for Option B, even in the case of $N=1$, the input context length remains no larger than that of standard digit-by-digit representations. We hypothesize that the $N=1$ configuration is optimal for Small Language Models (SLMs). In this setup, the representation is extremely compact (e.g., only 330 additional tokens to cover a range up to $10^{33}$), and the bijective nature of the tokens eliminates “semantic blur”. By providing an explicit one-to-one mapping between each digit and its magnitude, $N=1$ reduces the representational burden on constrained weight matrices, potentially leading to more robust arithmetic grounding in sub-billion parameter architectures.
@@ 6.2 Limitations and Future Work
  • Alternative grouping sizes: While the 3-digit grouping (base-1000) is natural for human readability, the framework allows for any $N$. For instance, Option B with $N=1$ might be more effective for SLMs, but other sizes could also be considered. Language-specific grouping (e.g., 3-2-2 in India, 4-4 in China) can be handled by the parser without changing the core vocabulary [12]. The final decision should be guided by which representation is most efficient for the specific model architecture.